To use 100 Calories of energy, how high would an 80.0 kg person have to climb?
1
Expert's answer
2020-10-23T11:53:04-0400

The energy, required to climb to some height, is

"E = mgh"

where g is gravitational acceleration.

We can convert given calories to metric units, namely

"E = 100 \; \textrm{cal} = 100 \cdot 4.18 \; J= 418 \;J".

So,

"\displaystyle h = \frac{E}{mg} = \frac{418}{80 \cdot9.81} = 0.53 \, m."


Answer: 0.53 m